OpenStudy (anonymous):

Find the perimeter of a rectangle with l = 1.2 m and w = 0.6 m. A. 0.72 m B. 1.8 m C. 3.6 m D. 7.2 m

OpenStudy (anonymous):

length=1.2 width=0.6

OpenStudy (anonymous):

add all of the sides, rectangle has 2sides the same and other 2 sides

OpenStudy (anonymous):

2*length+2*width

OpenStudy (nurali):

P= 2 * L + 2 * W P=2*1.2+2*0.6 P=2.4+1.2 P=________

OpenStudy (anonymous):

thank yall!!!

OpenStudy (nurali):

My Pleasure @kels200105

OpenStudy (anonymous):

:D